# On algebraic properties of some non-prime ideals of collections of cells

### v2

@@ Abstract.
In this work, we consider a family of coordinate rings obtained as a quotient ring of a binomial ideal associated to a collection of cells. The algebraic properties of Cohen-Macaulayness, Gorensteiness and the Hilbert-Poincaré series have been studied in literature only for some classes of collections of cells whose related binomial ideals are prime. We study these algebraic properties for some classes of collections of cells whose coordinate rings are not domain, considering a new strategy involving the combinatorics of the collections of cells and a recent result of Conca and Varbaro on the radical initial ideals.
@@ Introduction
Combinatorial Commutative Algebra has been an intensive area of research since the pioneering work of R. Stanley ([34]). In 2012 a new topic emerged in this field by the work of A. A. Qureshi. In [28] she establishes a connection between collections of cells and Commutative Algebra, assigning to every collection of cells $\mathcal{P}$ the ideal of the inner 2-minors of $\mathcal{P}$, in a suitable polynomial ring $S_{\mathcal{P}}$. This ideal $I_{\mathcal{P}}$ is called the inner 2-minor ideal of $\mathcal{P}$ and $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$ is said the coordinate ring of $\mathcal{P}$. In particular, if $\mathcal{P}$ is a polyomino, that is a collection of cells where the cells are joined edge by edge, then $I_{\mathcal{P}}$ is called the polyomino ideal of $\mathcal{P}$. The aim of the research consists in the studying of the main algebraic properties of $K[\mathcal{P}]$ depending on the shape of $\mathcal{P}$. Since 2012, many algebraic properties of the polyomino ideals have been studied and many open questions have arisen. For example, one of the most interesting is the classification of the polyominoes whose coordinate ring is a domain. As usual, we say that a collection of cells is prime if the related ideal is prime. In [18, 19, 30], it is proved that if $\mathcal{P}$ is simple, which means $\mathcal{P}$ has no hole, then $K[\mathcal{P}]$ is a domain. Nowadays, the study is devoted to non-simple polyominoes and, although there are several articles where it is discussed, a complete characterization is still unknown. We invite the reader to consult [3, 5, 18, 19, 21, 25, 26, 30, 33]. Also, many other algebraic properties, like when $K[\mathcal{P}]$ is Cohen-Macaulay or Gorenstein, are known only for some specific polyominoes and they are discussed just for prime classes of collections of cells so far. Regarding the Cohen-Macaulayness, the demonstrative strategy is based on two well-known theorems of Sturmfels ([36, Chapter 13]) and Hochster ([2, Theorem 6.3.5]), which state that a toric ring whose defining ideal admits a squarefree initial ideal is a normal Cohen-Macaulay ring. See for instance [6, 21, 26, 30]. No class of collections of cells or polyominoes, whose coordinate ring is a Cohen-Macaulay ring but not a domain, has been discussed. Motivated by these considerations, we investigate the Cohen-Macaulayness and the Hilbert-Poincaré series of the zig-zag collections, a class of non-prime collections of cells coming from the so-called zig-zag walks (see [25, Definition 3.2]), and of the closed paths, a kind of non-simple polyominoes introduced in [3]. Taking advantage of some recent results shown in [9], we provide a demonstrative technique involving the combinatorics of the collections of cells, like the combinatorial interpretation of the Krull dimension and the rook polynomial of a polyomino. The rook polynomial is a polynomial whose $k$-th coefficient is the number of ways of placing $k$ non-attacking rooks in a polyomino. The maximum number of non-attacking rooks is called the rook number of $\mathcal{P}$. The connection of the Castelnuovo-Mumford regularity of $K[\mathcal{P}]$ with the rook number is shown in [14] for the class of $L$-convex polyominoes. In general, the study of the $h$-polynomial in terms of the rook polynomial is a very interesting issue and this research has produced a relevant number of articles ([7, 11, 22, 24, 29, 31]). The rook polynomial also plays a crucial role in the study of the Gorensteiness of $K[\mathcal{P}]$ when it is a domain. A well-known result of Stanley ([35]) allows us to characterize the Gorenstein coordinate rings which are domain, checking if the $h$-polynomial is palindromic, which means counting the number of suitable configurations of non-attacking rooks. It is hard to characterize non-prime polyominoes whose coordinate ring is Gorenstein and such a class has never appeared so far. Some partial results are discussed in our paper (look at Corollaries 2.5, 3.9 and Theorem 4.18) and we point out that an approach using the initial ideal is not useful (Remark 3.10). Finally, for other very interesting combinatorial problems on polyomino ideals, we invite the reader to look at [1, 8, 10, 13, 15, 23, 32].
@@
The article is structured as follows. Section 1 is devoted to provide some basic definitions and properties of collections of cells and the associated coordinate ring and to recall some useful known results of Commutative Algebra. In Section 2 we prove that the $h$-polynomial of a simple collection of cells is equal to the product of the $h$-polynomial of each connected component (Theorem 2.2), exploiting the monomial order in [27] and [5, Theorem 3.3]. As a direct consequence, we obtain in Corollary 2.5 that the $h$-polynomial of a simple thin collection of cells coincides with the rook-polynomial. In Section 3 we introduce a new class of non-prime collections of cells, named zig-zag collections. We show that their coordinate rings are Cohen-Macaulay (Theorem 3.6), we compute the Hilbert-Poincaré series in terms of the switching rook polynomial (Corollary 3.8) and we eventually discuss the Gorenstein property (Corollary  3.9). The results achieved in these sections are used in Section 4 to prove our main result, which states that the coordinate ring of a non-prime closed path is Cohen-Macaulay (Theorem 4.13). In the end, Theorem 4.18 summarizes the results about the primality and the Cohen-Macaulayness of this kind of polyominoes and provides a complete characterization of the Gorenstein property in terms of the length of the maximal blocks. We conclude the paper with Section 5, giving some open questions for future works.
1. Collections of cells and known results on graded $K$-algebras
@@
Let $(i,j),(k,l)\in{\mathbb{Z}}^{2}$. We say that $(i,j)\leq(k,l)$ if $i\leq k$ and $j\leq l$. Consider $a=(i,j)$ and $b=(k,l)$ in ${\mathbb{Z}}^{2}$ with $a\leq b$. The set $[a,b]=\{(m,n)\in{\mathbb{Z}}^{2}:i\leq m\leq k,\ j\leq n\leq l\}$ is called an interval of ${\mathbb{Z}}^{2}$. If $i<k$ and $j<l$, then we say that $[a,b]$ is a proper interval. In this case, we call $a$ and $b$ the diagonal corners of $[a,b]$, and $c=(i,l)$ and $d=(k,j)$ the anti-diagonal corners of $[a,b]$. If $j=l$ (or $i=k$), then $a$ and $b$ are in horizontal (or vertical) position. A proper interval $C=[a,b]$ with $b=a+(1,1)$ is called a cell of ${\mathbb{Z}}^{2}$; moreover, the elements $a$, $b$, $c$ and $d$ are called respectively the lower left, upper right, upper left and lower right corners of $C$. The set of vertices of $C$ is $V(C)=\{a,b,c,d\}$ and the set of edges of $C$ is $E(C)=\{\{a,c\},\{c,b\},\{b,d\},\{a,d\}\}$. Let $\mathcal{S}$ be a non-empty collection of cells in $\mathbb{Z}^{2}$. Then $V(\mathcal{S})=\bigcup_{C\in\mathcal{S}}V(C)$ and $E(\mathcal{S})=\bigcup_{C\in\mathcal{S}}E(C)$, while the rank of $\mathcal{S}$ is the number of cells that belong to $\mathcal{S}$. If $C$ and $D$ are two distinct cells of $\mathcal{S}$, then a walk from $C$ to $D$ in $\mathcal{S}$ is a sequence $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of cells of ${\mathbb{Z}}^{2}$ such that $C_{i}\cap C_{i+1}$ is an edge of $C_{i}$ and $C_{i+1}$ for $i=1,\dots,m-1$. Moreover, if $C_{i}\neq C_{j}$ for all $i\neq j$, then $\mathcal{C}$ is called a path from $C$ to $D$.
We say that $C$ and $D$ are connected in $\mathcal{S}$ if there exists a path of cells in $\mathcal{S}$ from $C$ to $D$. A polyomino $\mathcal{P}$ is a non-empty, finite collection of cells in $\mathbb{Z}^{2}$ where any two cells of $\mathcal{P}$ are connected in $\mathcal{P}$ (see Figure 1 (a)) Let $\mathcal{P}$ be a non-empty, finite collection of cells in $\mathbb{Z}^{2}$, $\mathcal{P}$ is called weakly connected if for any two cells $C$ and $D$ in $\mathcal{P}$ there exists a sequence of cells $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of $\mathcal{P}$ such that $V(C_{i})\cap V(C_{i+1})\neq\emptyset$ for all $i=1,\dots,m-1$. A subset of cells $\mathcal{P}^{\prime}$ of $\mathcal{P}$ is called a connected component of $\mathcal{P}$ if $\mathcal{P}^{\prime}$ is a polyomino and it is maximal with respect to the set inclusion, that is if $A\in\mathcal{P}\setminus\mathcal{P}^{\prime}$ then $\mathcal{P}^{\prime}\cup\{A\}$ is not a polyomino. For instance, see Figure 1 (b). Observe trivially that every polyomino is a weakly connected collection of cells.
(a) A polyomino.
(b) A collection of cells.
Figure 1. The collection of cells in the right has three connected components.
@@
A sub-polyomino of $\mathcal{P}$ is a polyomino whose cells belong to $\mathcal{P}$. A weakly connected collection of cells is said to be thin if it does not contain the square tetromino (that is, the square obtained as a union of four distinct cells). We say that $\mathcal{P}$ is simple if for any two cells $C$ and $D$ not in $\mathcal{P}$ there exists a path of cells not in $\mathcal{P}$ from $C$ to $D$. A finite collection of cells $\mathcal{H}$ not in $\mathcal{P}$ is an hole of $\mathcal{P}$ if any two cells of $\mathcal{H}$ are connected in $\mathcal{H}$ and $\mathcal{H}$ is maximal with respect to set inclusion. For example, the collections of cells in Figure 1 are not simple. Each hole of $\mathcal{P}$ is a simple polyomino, and $\mathcal{P}$ is simple if and only if it has no hole. An interval $[a,b]$ with $a=(i,j)$, $b=(k,j)$ and $i<k$ is called a horizontal edge interval of $\mathcal{P}$ if the sets $\{(\ell,j),(\ell+1,j)\}$ are edges of cells of $\mathcal{P}$ for all $\ell=i,\dots,k-1$. In addition, if $\{(i-1,j),(i,j)\}$ and $\{(k,j),(k+1,j)\}$ do not belong to $E(\mathcal{P})$, then $[a,b]$ is called a maximal horizontal edge interval of $\mathcal{P}$. We define similarly a vertical edge interval and a maximal vertical edge interval.
@@
Figure 2. An example of a zig-zag walk of $\mathcal{P}$.
@@
Let $\mathcal{P}$ be a collection of cells. Two rooks in $\mathcal{P}$ are in non-attacking position if they do not belong to the same vertical or horizontal cell interval contained in $\mathcal{P}$. A $k$-rook configuration in $\mathcal{P}$ is a configuration of $k$ rooks which are arranged in $\mathcal{P}$ in non-attacking positions. Figure 3 shows a 6-rook configuration.
Figure 3. An example of a $6$-rook configuration in $\mathcal{P}$.
@@
The rook number $r(\mathcal{P})$ is the maximum number of rooks that can be arranged in $\mathcal{P}$ in non-attacking positions. For all $k\in\{0,\ldots,r(\mathcal{P})\}$, we denote by $\mathcal{R}_{k}$ the set of all $k$-rook configurations in $\mathcal{P}$, setting conventionally $R_{0}=\{\emptyset\}$. We also set $r_{k}=|\mathcal{R}_{k}|$ for all $k\in\{0,\dots,r(\mathcal{P})\}$. The rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $r_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}r_{k}t^{k}$. Moreover, the set $\mathcal{R}=\mathcal{R}_{0}\cup\mathcal{R}_{1}\dots\cup\mathcal{R}_{r(\mathcal{P})}$ is a simplicial complex, called rook complex. We call a rectangle of $\mathcal{P}$ the set of all cells belonging to an inner interval of $\mathcal{P}$ and we say that two cells $A,B$, with left lower corners $(i,j),(h,k)$ respectively, are in diagonal position if $(i+1,j+1)\leq(h,k)$, while they are in anti-diagonal) if $(i+1,j+1)$ and $(h,k)$ are not comparable with respect to $\prec$. Two non-attacking rooks in $\mathcal{P}$ are switching rooks if they are placed in diagonal or anti-diagonal cells of a rectangle of $\mathcal{P}$. In such a case we say that the rooks are in a diagonal or anti-diagonal position, respectively. Fix $k\in\{0,\dots,r_{\mathcal{P}}\}$. Let $F\in\mathcal{R}_{k}$ and $R_{1}$ and $R_{2}$ be two switching rooks in $F$ in diagonal (resp. anti-diagonal) position. Let $R_{1}^{\prime}$ and $R_{2}^{\prime}$ be the rooks in anti-diagonal (resp. diagonal) cells in $R$. Then the set $\left(F\backslash\{R_{1},R_{2}\}\right)\cup\{R_{1}^{\prime},R_{2}^{\prime}\}$ belongs to $\mathcal{R}_{k}$. The operation of replacing $R_{1}$ and $R_{2}$ by $R_{1}^{\prime}$ and $R_{2}^{\prime}$ is called switch of $R_{1}$ and $R_{2}$. This induce the following equivalence relation $\sim$ on $\mathcal{R}_{k}$: let $F_{1},F_{2}\in\mathcal{R}_{k}$, then $F_{1}\sim F_{2}$ if $F_{2}$ can be obtained from $F_{1}$ after some switches. We define the quotient set ~$\tilde{\mathcal{R}}_{k}=\mathcal{R}_{k}/\sim$. We set $\tilde{r}_{k}=|\tilde{\mathcal{R}}_{k}|$ for all $k\in[r(\mathcal{P})]$. The switching rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $\tilde{r}_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}\tilde{r}_{k}t^{k}$. Observe that if $\mathcal{P}$ is a thin polyomino then $\tilde{r}_{\mathcal{P}}(t)=r_{\mathcal{P}}(t).$ Moreover, let $\mathcal{P}$ be a collection of cells. Consider an element $\mathcal{C}$ of $\tilde{R}_{k}$ and we assume that all rooks in $\mathcal{C}$ are in diagonal position. We say that $\mathcal{C}$ is a canonical configuration in $\mathcal{P}$ of $k$ rooks, so $\tilde{r}_{k}$ is the number of canonical positions in $\mathcal{P}$ of $k$ rooks.
@@
Assume that $\mathcal{P}$ consists of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Denote by $\tilde{r}(t)$ the switching rook polynomial of $\mathcal{P}$ and we set
$\tilde{r}_{1}(t)=\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{i}\quad\text{and}\quad\tilde{r}_{2}(t)=\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}.$
Hence

### v3

# On Cohen-Macaulay non-prime collections of cells

@@ Abstract.
In this paper we investigate Cohen-Macaulayness, Gorensteinness and the Hilbert-Poincaré series for some classes of non-prime collections of cells. In particular, we show that all closed path polyominoes are Cohen-Macaulay and we characterize those that are Gorenstein.
@@ Introduction
Combinatorial Commutative Algebra has been an intensive area of research since the pioneering work of R. Stanley ([34]). In 2012 a new topic emerged in this field due to the work of A. A. Qureshi ([29]). She established a connection between collections of cells and Commutative Algebra, assigning to every collection $\mathcal{P}$ of cells the ideal of the inner 2-minors of $\mathcal{P}$ in a suitable polynomial ring $S_{\mathcal{P}}$. This ideal $I_{\mathcal{P}}$ is called the inner 2-minor ideal of $\mathcal{P}$ and $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$ is said to be the coordinate ring of $\mathcal{P}$. In particular, if $\mathcal{P}$ is a polyomino, that is a collection of cells where the cells are joined edge by edge, then $I_{\mathcal{P}}$ is called the polyomino ideal of $\mathcal{P}$. The aim of the research is to study the main algebraic properties of $K[\mathcal{P}]$ depending on the geometry of $\mathcal{P}$. The most investigated include primality, Cohen-Macaulayness and Gorensteinness. As usual, we say that a collection of cells is prime (respectively, Cohen-Macaulay or Gorenstein) if the related binomial ideal is prime (respectively, Cohen-Macaulay or Gorenstein). Regarding the classification of the primality for polyominoes, it is proven in [19, 20, 31] that if $\mathcal{P}$ is simple, which means that $\mathcal{P}$ has no hole, then $K[\mathcal{P}]$ is a normal Cohen-Macaulay domain. Nowadays, the study is devoted to non-simple polyominoes but a complete characterization of primality, Cohen-Macaulay and Gorenstein properties is still unknown, despite the efforts of many mathematicians (see [2, 4, 6, 8, 19, 20, 22, 26, 27, 30, 31, 32, 33]). Other very interesting combinatorial problems on polyomino ideals have been investigated as shown in [9, 11, 14, 16, 24].
Cohen-Macaulayness and Goresteiness have so far been discussed only for prime collections of cells. In all the papers where Cohen-Macaulayness is studied ([7, 22, 25, 27, 31]), the demonstrative strategy is based on two well-known theorems of Sturmfels ([36, Chapter 13]) and Hochster ([3, Theorem 6.3.5]), which state that a toric ring whose defining ideal admits a squarefree initial ideal is a normal Cohen-Macaulay ring. For Goresteiness, a well-known result of Stanley ([35]) allows to characterize the prime polyominoes which are Gorenstein, by checking if the $h$-polynomial is palindromic. In such a context, the so-called rook polynomial plays a crucial role. The rook polynomial is a polynomial whose $k$-th coefficient is the number of ways of placing $k$ non-attacking rooks in a polyomino. The maximum number of non-attacking rooks is called the rook number of $\mathcal{P}$. The connection of the Castelnuovo-Mumford regularity of $K[\mathcal{P}]$ with the rook number was shown in [15] for the first time. So far, the study of the $h$-polynomial in terms of the rook polynomial has produced a relevant number of articles, including [8, 12, 23, 25, 30, 32], and it could help to characterize the prime polyominoes that are Gorenstein. Indeed, if the $h$-polyonomial of a prime polyomino $\mathcal{P}$ is the rook polynomial, then we can check the palindromicity counting the number of suitable configurations of non-attacking rooks in $\mathcal{P}$. On the other hand, discussing non-prime polyominoes which are Gorenstein seems very hard and such classes have not appeared so far.
In this context a natural question arises: what can we say about the Cohen-Macaulay and the Gorenstein properties for non-prime collections of cells? Motivated by this question, we begin to investigate these properties for non-prime collections of cells. From [26, Corollary 3.6] we know that if a collection of cells contains the so-called zig-zag walks (see [26, Definition 3.2]), then the related binomial ideal is not prime. This suggests to consider as a first step a class of non-prime collections of cells made up of zig-zag walks, which we called zig-zag collections. As conjectured in [26, Conjecture 4.6], every non-prime collection of cells should contain a zig-zag collection, so the latter could be useful for gaining information on non-prime collections of cells, by applying a decomposition of them into suitable zig-zag collections. This method is used to study the before mentioned properties for the closed paths, a kind of non-simple polyominoes introduced for the first time in [4], in the non-prime case. In particular, for this class of polyominoes, we state the following result at the end of this paper:
Theorem 4.18 Let $\mathcal{P}$ be a closed path polyomino. Then:
  1. (1)
  2. (2)
    The $h$-polyonomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
  3. (3)
@@
The article is structured as follows. Section 1 is devoted to provide some basic definitions and properties of collections of cells and the associated coordinate rings. In Section 2 we prove that the $h$-polynomial of a simple collection of cells is equal to the product of the $h$-polynomial of each connected component (Theorem 2.2), exploiting the monomial order in [28] and [6, Theorem 3.3]. As a direct consequence, we can generalize [26, Theorem 3.13], obtaining in Corollary 2.5 that the $h$-polynomial of a simple thin collection of cells coincides with the rook-polynomial. In Section 3 we introduce a new class of non-prime collections of cells, named zig-zag collections. Taking advantage of some recent results shown in [10], we show that their coordinate rings are Cohen-Macaulay (Theorem 3.7), we compute the Hilbert-Poincaré series in terms of the switching rook polynomial (Corollary 3.8) and we eventually discuss the Gorenstein property (Corollary  3.9). The results achieved in the previous sections are collected in Section 4 to prove Theorem 4.18, which is our main result.
1. Collections of cells and their associated coordinate rings
@@
Let $(i,j),(k,l)\in{\mathbb{Z}}^{2}$. We say that $(i,j)\leq(k,l)$ if $i\leq k$ and $j\leq l$. Consider $a=(i,j)$ and $b=(k,l)$ in ${\mathbb{Z}}^{2}$ with $a\leq b$. The set $[a,b]=\{(m,n)\in{\mathbb{Z}}^{2}:i\leq m\leq k,\ j\leq n\leq l\}$ is called an interval of ${\mathbb{Z}}^{2}$. If $i<k$ and $j<l$, then we say that $[a,b]$ is a proper interval. In this case, we call $a$ and $b$ the diagonal corners of $[a,b]$, and $c=(i,l)$ and $d=(k,j)$ the anti-diagonal corners of $[a,b]$. If $j=l$ (or $i=k$), then $a$ and $b$ are in horizontal (or vertical) position. A proper interval $C=[a,b]$ with $b=a+(1,1)$ is called a cell of ${\mathbb{Z}}^{2}$; moreover, the elements $a$, $b$, $c$ and $d$ are called respectively the lower left, upper right, upper left and lower right corners of $C$. The set of vertices of $C$ is $V(C)=\{a,b,c,d\}$ and the set of edges of $C$ is $E(C)=\{\{a,c\},\{c,b\},\{b,d\},\{a,d\}\}$. Let $\mathcal{S}$ be a non-empty collection of cells in $\mathbb{Z}^{2}$. Then $V(\mathcal{S})=\bigcup_{C\in\mathcal{S}}V(C)$ and $E(\mathcal{S})=\bigcup_{C\in\mathcal{S}}E(C)$, while the rank of $\mathcal{S}$ is the number of cells that belong to $\mathcal{S}$. If $C$ and $D$ are two distinct cells of $\mathcal{S}$, then a walk from $C$ to $D$ in $\mathcal{S}$ is a sequence $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of cells of ${\mathbb{Z}}^{2}$ such that $C_{i}\cap C_{i+1}$ is an edge of $C_{i}$ and $C_{i+1}$ for $i=1,\dots,m-1$. Moreover, if $C_{i}\neq C_{j}$ for all $i\neq j$, then $\mathcal{C}$ is called a path from $C$ to $D$. We say that $C$ and $D$ are connected in $\mathcal{S}$ if there exists a path of cells in $\mathcal{S}$ from $C$ to $D$. A polyomino $\mathcal{P}$ is a non-empty, finite collection of cells in $\mathbb{Z}^{2}$ where any two cells of $\mathcal{P}$ are connected in $\mathcal{P}$ (see Figure 1 (a)) Let $\mathcal{P}$ be a non-empty, finite collection of cells in $\mathbb{Z}^{2}$, $\mathcal{P}$ is called weakly connected if for any two cells $C$ and $D$ in $\mathcal{P}$ there exists a sequence of cells $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of $\mathcal{P}$ such that $V(C_{i})\cap V(C_{i+1})\neq\emptyset$ for all $i=1,\dots,m-1$. A subset of cells $\mathcal{P}^{\prime}$ of $\mathcal{P}$ is called a connected component of $\mathcal{P}$ if $\mathcal{P}^{\prime}$ is a polyomino and it is maximal with respect to the set inclusion, that is, if $A\in\mathcal{P}\setminus\mathcal{P}^{\prime}$ then $\mathcal{P}^{\prime}\cup\{A\}$ is not a polyomino. For instance, see Figure 1 (b). Observe trivially that every polyomino is a weakly connected collection of cells.
Figure 1. A non-simple polyomino, a non-simple collection of cells, a simple thin polyomino and a simple thin collection of cells.
@@
A sub-polyomino of $\mathcal{P}$ is a polyomino whose cells belong to $\mathcal{P}$. A weakly connected collection of cells is said to be thin if it does not contain the square tetromino (that is, the square obtained as a union of four distinct cells); see Figures 1c and 1d. We say that $\mathcal{P}$ is simple if for any two cells $C$ and $D$ not in $\mathcal{P}$ there exists a path of cells not in $\mathcal{P}$ from $C$ to $D$. A finite collection of cells $\mathcal{H}$ not in $\mathcal{P}$ is an hole of $\mathcal{P}$ if any two cells of $\mathcal{H}$ are connected in $\mathcal{H}$ and $\mathcal{H}$ is maximal with respect to set inclusion. For example, the collections of cells in Figures 1a, 1b and 2 are not simple. Each hole of $\mathcal{P}$ is a simple polyomino, and $\mathcal{P}$ is simple if and only if it has no hole. An interval $[a,b]$ with $a=(i,j)$, $b=(k,j)$ and $i<k$ is called a horizontal edge interval of $\mathcal{P}$ if the sets $\{(\ell,j),(\ell+1,j)\}$ are edges of cells of $\mathcal{P}$ for all $\ell=i,\dots,k-1$. In addition, if $\{(i-1,j),(i,j)\}$ and $\{(k,j),(k+1,j)\}$ do not belong to $E(\mathcal{P})$, then $[a,b]$ is called a maximal horizontal edge interval of $\mathcal{P}$. We define similarly a vertical edge interval and a maximal vertical edge interval.
@@
(a) An example of a zig-zag walk of $\mathcal{P}$.
(b) An example of a $6$-rook configuration in $\mathcal{P}$.
Figure 2. Non-simple polyominoes.
See Figure 2a for an example.
@@
Let $\mathcal{P}$ be a collection of cells. Two rooks in $\mathcal{P}$ are in non-attacking position if they do not belong to the same vertical or horizontal cell interval contained in $\mathcal{P}$. A $k$-rook configuration in $\mathcal{P}$ is a configuration of $k$ rooks which are arranged in $\mathcal{P}$ in non-attacking positions. Figure 2b shows a 6-rook configuration.
@@
The rook number $r(\mathcal{P})$ is the maximum number of rooks that can be arranged in $\mathcal{P}$ in non-attacking positions. For all $k\in\{0,\ldots,r(\mathcal{P})\}$, we denote by $\mathcal{R}_{k}$ the set of all $k$-rook configurations in $\mathcal{P}$, using the conventional setting $R_{0}=\{\emptyset\}$. We also set $r_{k}=|\mathcal{R}_{k}|$ for all $k\in\{0,\dots,r(\mathcal{P})\}$. The rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $r_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}r_{k}t^{k}$. Moreover, the set $\mathcal{R}=\mathcal{R}_{0}\cup\mathcal{R}_{1}\dots\cup\mathcal{R}_{r(\mathcal{P})}$ is a simplicial complex, called rook complex. We call a rectangle of $\mathcal{P}$ the set of all cells belonging to an inner interval of $\mathcal{P}$ and we say that two cells $A,B$, with left lower corners $(i,j),(h,k)$ respectively, are in diagonal position if $(i+1,j+1)\leq(h,k)$, while they are in anti-diagonal) if $(i+1,j+1)$ and $(h,k)$ are not comparable with respect to $\prec$. Two non-attacking rooks in $\mathcal{P}$ are switching rooks if they are placed in diagonal or anti-diagonal cells of a rectangle of $\mathcal{P}$. In such a case we say that the rooks are in a diagonal or anti-diagonal position, respectively. Fix $k\in\{0,\dots,r_{\mathcal{P}}\}$. Let $F\in\mathcal{R}_{k}$ and $R_{1}$ and $R_{2}$ be two switching rooks in $F$ in diagonal (resp. anti-diagonal) position. Let $R_{1}^{\prime}$ and $R_{2}^{\prime}$ be the rooks in anti-diagonal (resp. diagonal) cells in $R$. Then the set $\left(F\backslash\{R_{1},R_{2}\}\right)\cup\{R_{1}^{\prime},R_{2}^{\prime}\}$ belongs to $\mathcal{R}_{k}$. The operation of replacing $R_{1}$ and $R_{2}$ by $R_{1}^{\prime}$ and $R_{2}^{\prime}$ is called switch of $R_{1}$ and $R_{2}$. This induces the following equivalence relation $\sim$ on $\mathcal{R}_{k}$: let $F_{1},F_{2}\in\mathcal{R}_{k}$, then $F_{1}\sim F_{2}$ if $F_{2}$ can be obtained from $F_{1}$ after some switches. We define the quotient set ~$\tilde{\mathcal{R}}_{k}=\mathcal{R}_{k}/\sim$. We set $\tilde{r}_{k}=|\tilde{\mathcal{R}}_{k}|$ for all $k\in[r(\mathcal{P})]$. The switching rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $\tilde{r}_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}\tilde{r}_{k}t^{k}$. Observe that if $\mathcal{P}$ is a thin polyomino then $\tilde{r}_{\mathcal{P}}(t)=r_{\mathcal{P}}(t).$ Moreover, let $\mathcal{P}$ be a collection of cells. Consider an element $\mathcal{C}$ of $\tilde{R}_{k}$ and we assume that all rooks in $\mathcal{C}$ are in diagonal position. We say that $\mathcal{C}$ is a canonical configuration in $\mathcal{P}$ of $k$ rooks, so $\tilde{r}_{k}$ is the number of canonical positions in $\mathcal{P}$ of $k$ rooks.
@@
Assume that $\mathcal{P}$ consists of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Denote by $\tilde{r}(t)$ the switching rook polynomial of $\mathcal{P}$ and we set $\tilde{r}_{1}(t)=\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{i}$ and $\tilde{r}_{2}(t)=\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}.$ Hence